### v3

# STAvatar: Soft Binding and Temporal Density Control for Monocular 3D Head Avatars Reconstruction

Abstract
Reconstructing high-fidelity and animatable 3D head avatars from monocular videos remains a challenging yet essential task. Existing methods based on 3D Gaussian Splatting typically bind Gaussians to mesh triangles and model deformations solely via Linear Blend Skinning, which results in rigid motion and limited expressiveness. Moreover, they lack specialized strategies to handle frequently occluded regions (e.g., mouth interiors, eyelids). To address these limitations, we propose STAvatar, which consists of two key components: (1) a UV-Adaptive Soft Binding framework that leverages both image-based and geometric priors to learn per-Gaussian feature offsets within the UV space. This UV representation supports dynamic resampling, ensuring full compatibility with Adaptive Density Control (ADC) and enhanced adaptability to shape and textural variations. (2) a Temporal ADC strategy, which first clusters structurally similar frames to facilitate more targeted computation of the densification criterion. It further introduces a novel fused perceptual error as clone criterion to jointly capture geometric and textural discrepancies, encouraging densification in regions requiring finer details. Extensive experiments on four benchmark datasets demonstrate that STAvatar achieves state-of-the-art reconstruction performance, especially in capturing fine-grained details and reconstructing frequently occluded regions. Our project is available at https://jiankuozhao.github.io/STAvatar/.
1 Introduction
Reconstructing animatable and photo-realistic 3D head avatars has been a longstanding problem in computer vision and graphics [3, 22, 29, 34, 26, 28], driven by growing demands in AR/VR, telepresence, digital humans and interactive media. To achieve high-quality avatar reconstruction, conventional methods typically rely on expensive multi-camera systems [7, 32] which require complex manual setups and are impractical for consumer-level deployment. In contrast, monocular-based approaches [25, 39, 6] aim to produce high-fidelity results using a single consumer-grade camera, enabling broader accessibility in immersive and personalized virtual environments.
Figure 1: STAvatar proposes a Soft Binding framework and a Temporal Adaptive Density Control strategy to reconstruct high-fidelity 3D head avatars from monocular videos.
Recent years have witnessed the emergence of 3D Gaussian Splatting (3DGS) [8] as a powerful volumetric representation, delivering both high-fidelity rendering and real-time performance in neural scene reconstruction. Building on its success in static scenes, many works have explored its application to animatable 3D head avatar reconstruction. These methods typically follow a two-stage pipeline: a parametric head-tracking algorithm [20, 44] is first employed to estimate the facial pose, expression, and shape for each frame; subsequently, Gaussian primitives are bound to mesh triangles and driven by a deformation field, which is typically implemented using Linear Blend Skinning (LBS) [19]. During training, each Gaussian’s parameters are optimized via differentiable rendering to minimize photometric loss. Despite achieving enhanced visual fidelity and animation realism, existing Gaussian-based approaches still face several open challenges. To address these limitations, we propose STAvatar, as shown in Fig. 1, and provide a detailed explanation below.
First, existing animatable reconstruction methods often suffer from rigid connection between Gaussian primitives and mesh triangles, resulting in poor handling of non‑rigid and fine‑scale deformations. Fig. 2(a) illustrates this simple hard binding framework [35, 24, 33] in which each Gaussian deforms solely according to the LBS of its parent triangle. Although this enables basic animation control, it fails to model fine-grained deformations, as the Gaussians remain relatively static within the triangle’s local coordinate frame. This limitation stems from the gap between mesh and Gaussian representations. While meshes inherently encode topology and support native skinning for deformation, Gaussians are structureless and rely on externally defined deformation fields to capture complex motions and fine details such as facial wrinkles. Some methods [1, 30, 5] attempt to enhance each Gaussian’s deformation using a fixed-dimension multilayer perceptron (MLP). However, this design requires predefining a fixed number of Gaussians, which is suboptimal since different identities often demand varying Gaussian densities. Besides, it neglects spatial context and may be incompatible with the Adaptive Density Control (ADC) inherent to 3DGS. To enhance deformation ability while preserving the flexibility of ADC, STAvatar proposes a UV-Adaptive Soft Binding framework, which integrates the LBS for coarse deformation and a dual-branch network for fine-detail recovery. By generating a feature offset map in UV space, STAvatar leverages spatial context to model the Gaussian deformation field and enables arbitrary-resolution sampling to support ADC.
Second, the vanilla ADC in 3DGS is tailored for static scene reconstruction and fails to handle transiently visible or frequently occluded regions in dynamic avatar reconstruction. As illustrated in Fig. 2(b), certain regions (e.g., mouth interiors) are only visible in a subset of frames and contribute minimally to rendering when occluded. This results in low average criterion-such as the positional gradient used in 3DGS-which impedes sufficient densification. In addition, as shown in Fig. 2(c), the positional gradient in 3DGS only captures geometric discrepancies while neglecting texture details [21], which leads to the omission of regions with high texture errors and the failure to add Gaussians to these regions. To address these limitations of ADC, STAvatar propose a Temporal Adaptive Density Control strategy incorporating FLAME-Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE-AP). As shown in Fig. 2(b), FTC partitions frames into structure‑coherent clusters to encourage frequently occluded regions remain visible within a cluster, facilitating targeted refinement. FPE‑AP replaces positional gradients by directly estimating each Gaussian’s fused perceptual error, which jointly considers both geometric and textural discrepancies. Besides, FPE-AP selects Gaussians exhibiting the highest instantaneous errors across iterations to better capture regions with large peak errors.
Figure 2: Limitations of existing research. (a) Hard binding forces Gaussians to remain relatively static within the triangle coordinate frames, thereby limiting their ability to capture fine-grained details. (b) Transiently visible regions, such as mouth interiors, often exhibit low average positional gradients, which impedes effective Gaussian densification. (c) The positional gradient only reflects geometric inconsistencies and often loses texture details, which hinders the addition of Gaussians in high-frequency regions.
Finally, we evaluate STAvatar through extensive experiments on four datasets, demonstrating clear improvements over existing methods in recovering fine details and handling challenging regions such as mouth interiors and eyelids. Besides, STAvatar exhibits accurate cross-reenactment performance, highlighting its potential in real-world applications. Our main contributions are as follows:
  • We propose a UV‑Adaptive Soft Binding framework that supports both soft binding and ADC of 3DGS in animatable Gaussian avatar reconstruction.
  • We propose a Temporal ADC strategy incorporating FTC and FPE-AP which enhances the ADC in dynamic Gaussian avatar reconstruction.
  • We conduct experiments on various datasets, demonstrating superior rendering quality, improved detail reconstruction, and more accurate cross-reenactment.
2 Related Work
2.1 Gaussian Head Reconstruction
3DGS [8] has recently emerged as a flexible and efficient approach for 3D scene representation. By employing a large number of anisotropic Gaussian primitives—each parameterized by geometry and appearance attributes—it achieves high-quality reconstructions, making it particularly well-suited for detailed human head modeling. To enable avatar driving, typical methods [19, 16, 25, 23, 24, 35] adopt a hard-binding framework that associates Gaussians with mesh triangles to transfer mesh motion via LBS, which, however, fails to bridge the inherent gap between mesh and Gaussian representations, leading to suboptimal performance in handling complex motions and expressions. Alternatively, other approaches [30, 1, 31, 14] leverage MLP to predict per-Gaussian attribute offsets, thereby achieving a more flexible soft-binding framework. However, these methods often neglect ADC by using a fixed number of Gaussians across different identities, which limits expressiveness and reconstruction quality. In addition, several studies [17, 12, 43] propose to represent avatars with a base model and a set of expression blendshapes. Although these methods can achieve high training efficiency, they often suffer from low fidelity or produce blurred results under large pose variations, since the learned blendshapes may not adequately capture diverse pose-dependent deformations.
2.2 Adaptive Density Control of 3DGS
ADC strategy is a key module in 3DGS which dynamically clones, splits and prunes Gaussians to balance detail preservation and representation compactness. The vanilla ADC selects primitives for densification based on average positional gradients and Gaussian scales. However, this heuristic often results in underfitting and blurred details, as it neglects texture-related cues and only focus average criterion. To overcome these limitations, some recent works [21, 37, 10, 18, 41] revise the densification criterion to incorporate richer information, thereby achieving finer details in underrepresented regions. Besides, optimization-theoretic approaches have been proposed to establish more principled rules for cloning and splitting: SteepGS [27] leverages steepest-descent analysis to trigger Gaussian splits, whereas 3DGS-MCMC [9] formulates density control as a Markov Chain Monte Carlo (MCMC) sampling process with $L_{1}$-regularized pruning. More recently, Efficient Density Control [2] introduces long-axis splitting operations and recovery-aware pruning strategies to reduce redundant Gaussians and accelerate convergence, while Gradient-Direction-Aware ADC [42] incorporates gradient coherence and angular alignment to more effectively determine which Gaussians should be cloned or split. However, as these methods are primarily tailored for static scene reconstruction, they tend to under-fit transiently visible regions and fail to jointly consider both geometric and textural discrepancies.
Figure 3: Overview of STAvatar. (a) In addition to a fixed identity reference image and its UV position map, we further rasterize the vertex offsets between reference mesh and control mesh to obtain a UV displacement map as input. (b) We construct a dual-branch network to predict a feature offset map in UV space, from which an offset $\delta_{i}$ is sampled for each Gaussian $g_{i}$. This offset is added to the coarsely estimated parameters $\tilde{\theta}$ to get final parameters $\theta^{*}$. The final images are then rendered using Gaussian Splatting. (c) We first construct a perceptual error map by combining $\mathcal{L}_{1}$ map and $\mathcal{L}_{\mathrm{d\text{-}ssim}}$ map. Then, we estimate the 2D projection of each Gaussian $g_{i}$ using the recorded attributes, based on which the fused perceptual error is computed.
3 Method
We first introduce the general Gaussian avatar reconstruction pipeline in Sec. 3.1. Next, we present a UV-Adaptive Soft Binding framework in Sec. 3.2. The Temporal Adaptive Density Control strategy, which integrates FLAME‑Conditioned Temporal Clustering (FTC) and a Fused Perceptual Error with Average-Peak Criterion (FPE‑AP), is discussed in Sec. 3.3. Finally, the detailed optimization process is described in Sec. 3.4.
3.1 Preliminaries
Initialization. Following [19], we initialize Gaussian primitives $g_{i}$ based on the tracked FLAME model. Each $g_{i}$ is bound to a mesh triangle and parameterized by a center position $\bm{\mu}$, a scale matrix $\bm{S}$, and a rotation matrix $\bm{R}$ defined on the canonical FLAME mesh [13]. The 3D Gaussian density is defined as:
$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\Sigma^{-1}(\bm{x}-\bm{\mu})},$ (1)
where $\Sigma=\bm{R}\,\bm{S}\,\bm{S^{\top}}\,\bm{R^{\top}}$ to keep the covariance matrix positive. In addition to geometric parameters, each primitive also carries appearance attributes: opacity $\alpha$ and color $c$.
Animate. To animate these canonical primitives according to FLAME deformations, we first transform their parameters using the parent triangle’s barycentric mapping to get the coarsely estimated parameters $\tilde{\theta}=\{\bm{\tilde{\mu}},\bm{\tilde{s}},\bm{\tilde{r}},\tilde{\alpha},\tilde{c}\}$. Specifically,
$\bm{\tilde{r}}=\bm{r}\bm{R},\ \bm{\tilde{\mu}}=k\bm{r}\bm{\mu}+\bm{t},\ \bm{\tilde{s}}=k\bm{s},\ \tilde{\alpha}=\alpha,\ \tilde{c}=c,$ (2)
where $\bm{r}$ and $\bm{t}$ denote the triangle’s relative rotation and barycenter translation, respectively, and the isotropic scale $k$ reflects the triangle’s size.
Render. Given the final parameters $\theta^{*}$ (see Sec. 3.2), the rendered color $\mathbf{C}$ of a pixel is obtained by compositing all overlapping primitives in a depth-sorted compositing:
$\bm{C}=\sum_{i=1}^{N}c_{i}^{*}\,\alpha_{i}^{\prime}\!\!\prod_{j=1}^{i-1}\!(1-\alpha_{j}^{\prime}),$ (3)
where $c_{i}^{*}$ is the RGB color encoded via 3rd-degree spherical harmonics, and $\alpha_{i}^{\prime}$ is the 2D projected opacity obtained by evaluating the Gaussian’s contribution in image space. This compositing accounts for occlusion by sorting the primitives according to their depth values.
3.2 UV-Adaptive Soft Binding Framework
Eq. (2) only yields a coarse deformation for each Gaussian $g_{i}$, failing to capture fine details or appearance changes since opacity $\tilde{\alpha}$ and color $\tilde{c}$ remain constant under LBS. To address this while preserving the flexibility of ADC, we introduce the UV‑Adaptive Soft Binding framework.
Input Preparation. As shown in Fig. 3(a), after tracking the flame model for video frames, we select a fixed reference image $Img_{r}$ to extract the texture input and rasterize its UV position map $UV_{pos}$ to obtain UV coordinate information. To further enhance the capacity for deformation modeling, we rasterize the vertex offsets between the reference mesh and the control mesh into UV space to generate an additional input $UV_{disp}$.
Soft Binding. Fig. 3(b) illustrates our dual-branch network, which estimates per-Gaussian feature offsets in UV space to capture rich spatial context. We first extract texture features $T$ by encoding $Img_{r}$ with a U-Net encoder, denoted as $\mathcal{E}_{i}$, such that $T=\mathcal{E}_{i}(Img_{r})$. These features are then concatenated with Fourier positional encoded position maps $UV_{pos}^{{}^{\prime}}=\mathcal{E}_{f}(UV_{pos})$ and displacement maps $UV_{disp}^{{}^{\prime}}=\mathcal{E}_{f}(UV_{disp})$. Besides, both branches $\Phi_{g}$ and $\Phi_{l}$ take as input a control code $\beta$, formed by concatenating expression, translation, and pose codes, to model Gaussian deformations. The decoded outputs, $\omega_{g}$ and $\omega_{l}$, are then fused to generate the final feature offset map $\Delta_{map}\in\mathbb{R}^{256\times 256\times 13}$, where each texel stores a 13‑dimensional Gaussian offset. The entire process can be described as:
$\omega_{g}=\Phi_{g}(T,UV_{pos}^{{}^{\prime}},\beta),$ (4)
$\omega_{l}=\sum_{i=1}^{4}H_{i}(M_{i}\odot\Phi_{l}(T,UV_{disp}^{{}^{\prime}},\beta)),$ (5)
$\Delta_{\!map}=\mathcal{F}(\omega_{g},\omega_{l}),$ (6)
where $M_{i}\in\{0,1\}^{256\times 256}$ denotes the UV spatial mask for the $i$-th facial region, including the eyes, mouth, nose, and forehead, and $\odot$ denotes the element-wise (Hadamard) product. The local branch $\phi_{l}$ shares a common decoder across all regions and applies region-specific decoding heads $H_{i}$ to the masked features. Note that $Img_{r}$ can be selected randomly; in our experiments, we default to using the first frame.
UV-Adaptive Sampling. We sample each Gaussian $g_{i}$’s offset $\delta_{i}=\{\delta_{\mu},\delta_{s},\delta_{r},\delta_{\alpha},\delta_{c}\}$ from the feature offset map $\Delta_{map}$ by assigning each $g_{i}$ a UV coordinate via our UV-Adaptive Sampling, which is updated upon densification to accommodate dynamic changes in Gaussian count. A brief algorithmic description of this sampling procedure is provided in Algorithm 1, with a detailed version included in the supplementary material. After sampling, we apply different activation functions to ensure numerical stability and constrain the effect of the predicted offsets $\delta_{i}$ (details in Appendix). The final Gaussian parameters $\theta^{*}$ are then computed as:
$\mu^{*}$ $=\tilde{\mu}+\delta_{\mu},\quad c^{*}=\tilde{c}+\delta_{c},\quad\alpha^{*}=\tilde{\alpha}+\delta_{\alpha},$ (7)
$s^{*}$ $=\tilde{s}\odot\delta_{s},\quad r^{*}=q(\tilde{r},\delta_{r}),$
where $q(\cdot,\cdot)$ represents the Hamilton product between two unit quaternions, used to compose the rotation $\tilde{r}$ and its predicted offset $\delta_{r}$. After obtaining the final parameters $\theta^{*}$, we employ Eq. (3) to render the images.
Algorithm 1 UV-Adaptive Sampling
Input: Face bindings $\mathcal{B}$, UV vertices $\mathcal{V}_{uv}$, UV faces $\mathcal{F}_{uv}$
Output: $\mathcal{U}\in\mathbb{R}^{N\times 2}$: UV coordinates per Gaussian
[…]
2: Build per-face pixel pool $\mathcal{P}$ from $\mathcal{M}_{face}$
[…]
4: Initialize empty barycentric buffer $\mathbf{b}$ and face index buffer $\mathbf{f}_{idx}$
5:for each face $f$ in face bindings $\mathcal{B}$ do
6:  $C_{f}\leftarrow$ count of points bound to face $f$
7:  $\mathcal{I}_{f}\leftarrow$ indices of points bound to face $f$
8:  if $f\in\mathcal{F}_{valid}$ then
9:   $\mathbf{p}_{f}\leftarrow$ pixel list for face $f$ from $\mathcal{P}$
10:   Sample $C_{f}$ pixels from $\mathbf{p}_{f}$
11:   Extract corresponding barycentric coordinates $\mathbf{b}_{f}$
12:  else
13:   Analytically sample $C_{f}$ barycentric coordinates $\mathbf{b}_{f}$
14:  end if
15:  Assign $\mathbf{b}_{f}$ to $\mathbf{b}[\mathcal{I}_{f}]$, and fill $\mathbf{f}_{idx}[\mathcal{I}_{f}]\leftarrow f$
16:end for
[…]
3.3 Temporal Adaptive Density Control
FPE‑AP. To identify Gaussians that require cloning, we introduce a more interpretable and comprehensive criterion, FPE-AP, which directly indicates regions exhibiting significant geometric or texture errors that may benefit from more Gaussian primitives. Since the error at each pixel is influenced by multiple overlapping Gaussians, it is computationally intractable and inefficient to compute the exact contribution of each individual Gaussian.
To address this, as illustrated in Fig. 3(c), we first construct a fused perceptual error map:
$E=(1-\lambda_{1})\,\left\lvert\mathcal{L}_{1}\right\rvert+\lambda_{1}\,\mathcal{L}_{\mathrm{d\text{-}ssim}},$ (8)
where $\mathcal{L}_{1}$ denotes the per-pixel absolute difference, and $\mathcal{L}_{\mathrm{d\text{-}ssim}}$ represents the per-pixel dissimilarity, derived from the Structural Similarity Index (SSIM). We set $\lambda_{1}=0.2$ to keep it consistent with the RGB loss design in Eq. (12), which better aligns with human perceptual sensitivity.
To estimate each Gaussian’s 2D projection for fused perceptual error calculation, we record its screen-space center $(x_{i},y_{i})$, total number of covered pixels $C_{i}$, and accumulated alpha blending weight $A_{i}=\sum_{p\in\mathcal{P}_{i}}a_{i}(p)$, where $a_{i}(p)$ is the blending weight of $g_{i}$ at pixel $p$. The influence region is defined as a square centered at $(x_{i},y_{i})$ with half-extent $R_{i}=\left\lfloor\frac{\sqrt{C_{i}}}{2}\right\rfloor$, and the corresponding pixel set $\mathcal{P}_{i}$ includes all pixels within this region. We then approximate the average fused perceptual error for $g_{i}$ as
$\bar{E}_{i}=\frac{A_{i}}{C_{i}}\sum_{p\in\mathcal{P}_{i}}E(p),$ (9)
where all window-sum computations are efficiently accelerated using a two-dimensional summed-area table (details in the supplementary material).
Furthermore, to capture transient spikes, we define a peak fused perceptual error across all iterations $t$ as
$E_{i}^{\mathrm{peak}}=\max_{t}\left(\frac{A_{i}^{(t)}}{C_{i}^{(t)}}\sum_{p\in\mathcal{P}_{i}^{(t)}}E^{(t)}(p)\right),$ (10)
and select top 3% Gaussians by $E_{i}^{\mathrm{peak}}$ to form the set $\mathcal{S}_{\mathrm{peak}}$. Finally, Gaussian $g_{i}$ is cloned if
$\bar{E}_{i}>\tau_{\mathrm{avg}}\quad\text{or}\quad i\in\mathcal{S}_{\mathrm{peak}},$ (11)
where $\tau_{\mathrm{avg}}=1\times 10^{-3}$. Note that, for Gaussian splitting, we still use the positional gradient as criterion since splitting is primarily driven by geometric inconsistency.
Figure 4: FLAME‑Conditioned Temporal Clustering. We cluster video frames into $K$ clusters and conduct ADC within each cluster’s training.
FTC. As shown in Fig. 4, we cluster video frames into $K$ clusters based on their FLAME parameters (expression, pose, and translation), with respective weights of 0.3, 0.6, and 0.1. Before clustering, we apply Principal Component Analysis (PCA) to reduce the parameter dimensionality for more effective distance computation. We then perform K-means clustering, which ensures the densification criterion is computed among structurally similar frames. This strategy mitigates insufficient densification in transiently visible regions. To accommodate different identities, we select $K$ within $[5,12]$ by maximizing the average silhouette score.
During training, each cluster is first optimized for NMN{-}M epochs, during which ADC is applied. After all clusters have been trained, we conduct additional $M$ epochs using randomly shuffled data to mitigate potential inconsistencies introduced by inter-cluster variations. In our experiments, we set $N=6$ and $M=1$.
Table 1: Comparison of quantitative results with state-of-the-art methods. The best and second are highlighted, respectively.
Method INSTA Dataset PointAvatar Dataset NerFace Dataset HDTF Dataset
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
SA [24] 27.48 0.9329 0.1046 24.93 0.8938 0.1478 26.14 0.9238 0.1145 26.02 0.8922 0.1821
MGA [1] 27.34 0.9351 0.0859 27.91 0.9277 0.1113 26.52 0.9333 0.0900 27.04 0.9171 0.1225
GA [19] 26.98 0.9378 0.0851 24.62 0.9073 0.1333 25.74 0.9279 0.0965 25.08 0.8879 0.1763
FA [30] 27.90 0.9357 0.0563 26.19 0.9055 0.0947 26.96 0.9331 0.0555 26.83 0.9110 0.0774
RGBA [12] 28.41 0.9493 0.0518 26.67 0.9217 0.0883 27.13 0.9438 0.0550 26.72 0.9277 0.0821
Fate [35] 28.33 0.9446 0.0508 28.36 0.9287 0.0776 27.12 0.9364 0.0558 27.18 0.9187 0.0849
Ours 30.63 0.9587 0.0304 28.25 0.9337 0.0495 30.08 0.9567 0.0311 27.99 0.9315 0.0542
Figure 5: Qualitative results of head avatar reconstruction. Our method recovers finer details and delicate structures such as wrinkles and teeth. Moreover, it produces clearer results in challenging regions like mouth interiors and eyelids.
3.4 Training Objective and Optimization
RGB Loss. We supervise the rendered images using a combination of $\mathcal{L}_{1}$ term and D-SSIM term following [8]. To better capture high-frequency facial details, we further introduce the perceptual loss $\mathcal{L}_{\mathrm{vgg}}$, which is activated only in the second half of training to balance performance and efficiency. The overall image loss is defined as:
$\mathcal{L}_{\mathrm{rgb}}=(1-\lambda_{1})\mathcal{L}_{1}+\lambda_{1}\mathcal{L}_{\mathrm{d\text{-}ssim}}+\gamma\lambda_{2}\mathcal{L}_{\mathrm{vgg}},$ (12)
where $\lambda_{1}=0.2$, $\lambda_{2}=0.05$, and $\gamma=1$ only during the latter half of training.
Regularization Loss. To regularize the predicted offset $\delta_{i}$, we apply a regularization term to constrain the scale and color offsets:
$\mathcal{L}_{\mathrm{offset}}=\lambda_{3}|\delta_{s}-1|+\lambda_{4}\delta_{c},$ (13)
where $\lambda_{3}=0.01$ and $\lambda_{4}=0.001$. Following GaussianAvatars [19], we also employ a position loss and a scale loss to regularize the spatial alignment and local scale of the 3D Gaussians with respect to their parent triangles.
The total training loss is thus defined as:
$\mathcal{L}=\mathcal{L}_{\mathrm{rgb}}+\mathcal{L}_{\mathrm{offset}}+\lambda_{p}\mathcal{L}_{\mathrm{position}}+\lambda_{s}\mathcal{L}_{\mathrm{scale}},$ (14)
where $\lambda_{p}=0.01$ and $\lambda_{s}=1$.
Optimization. We employ Adam [11] for parameter optimization, using identical hyperparameter settings across all subjects. The learning rate is set to $1\times 10^{-4}$ for the UV-Adaptive Soft Binding framework, while the learning rates for the remaining parameters follow those used in 3D Gaussian Splatting. In addition to the Gaussian Splatting parameters, we also finetune the expression, translation and joint rotation parameters of FLAME for each timestep as done in GaussianAvatars. As each Gaussian is bound to a mesh triangle, we do not observe noticeable floating Gaussians and thus skip opacity resetting in ADC.
4 Experiments
4.1 Experimental Setup
Figure 6: Qualitative results of cross-identity reenactment. Our method accurately animates the source avatar performing expressions such as smiling and eye-closing.
Implementation Details. We conduct both training and inference on a single NVIDIA RTX 3090 GPU. Each identity’s video is trained for 6 epochs. All videos are cropped and resized to a resolution of $512\times 512$. A robust matting method [15] is applied to extract the foreground with the background set to white. For avatar initialization, we adopt VHAP [20] for monocular head tracking.
Dataset. We evaluate our method on four datasets, including INSTA [45], PointAvatar [40], Nerface [4], and HDTF [38], comprising a total of 22 identities. For PointAvatar, we follow the original data split strategy in [40], while for the other datasets, the last 450 frames are reserved as the test set.
Baselines. We compare our method against six state-of-the-art Gaussian avatar reconstruction approaches: GaussianAvatars (GA) [19], FateAvatar (Fate) [35], RGAvatar (RGBA) [12], FlashAvatar (FA) [30], MonoGaussianAvatar (MGA) [1], and SplattingAvatar (SA) [24]. Note that simply equalizing or increasing training epochs across methods may not yield optimal results. Accordingly, we follow the settings in the original papers, setting 10 epochs for GA, Fate, and RGBA, 20 for FA, 30 for SA, and 100 for MGA to allow each method to converge properly. Since GA is originally designed for multi-view reconstruction, we adopt various settings from original papers for fair comparison.
4.2 Evaluation
Quantitative Results. We select PSNR, SSIM, and LPIPS [36] as evaluation metrics to comprehensively assess the visual quality of the reconstructed avatars. As shown in Tab. 1, our method consistently outperforms existing state-of-the-art approaches across most datasets and metrics, achieving notable improvements in reconstruction quality. In particular, our method attains the highest SSIM scores and the lowest LPIPS values on all four datasets, demonstrating its strong ability to preserve both geometric accuracy and perceptual fidelity.
Qualitative Results. As shown in Fig. 5, owing to the non-rigid and flexible deformation capability of our UV-Adaptive Soft Binding framework, our method effectively recovers fine-grained details and subtle structures, such as wrinkles and hair. Furthermore, benefiting from the Temporal ADC strategy, our method can produce clearer and higher-quality reconstructions in challenging regions like mouth interiors and eyelids, which are only visible in a subset of frames and are typically difficult to reconstruct. We further evaluate our method on cross-identity reenactment as a real-world test for avatar animation. As illustrated in Fig. 6, our method accurately animates the source avatar performing expressions such as smiling and eye-closing, while faithfully preserving the identity-specific characteristics and fine-grained details.
Table 2: Ablation quantitative results on the INSTA dataset.
Method PSNR↑ SSIM↑ LPIPS↓
w/o Soft Binding 29.66 0.9525 0.0398
w/o ADC 30.31 0.9529 0.0505
w/o FPE-AP 30.38 0.9574 0.0321
w/o FTC 30.43 0.9579 0.0310
w/o Local Decoder $\phi_{l}$ 30.36 0.9575 0.0306
w/o $\mathcal{L}_{\mathrm{vgg}}$ 30.58 0.9592 0.0426
Ours 30.63 0.9587 0.0304
Figure 7: Qualitative results of the ablation study on marcel case. Our full method produces more realistic textures and geometric structures in transiently visible regions, such as the mouth interior, and captures more fine details.
4.3 Ablation Study
We conduct an ablation study on self-reenactment using the INSTA dataset. The quantitative results are presented in Tab. 2, while the qualitative comparisons are shown in Fig. 7, demonstrating the effectiveness of each component. Additionally, we provide the ablation analysis of all hyper-parameters in the supplementary materials.
Figure 8: Comparison of Gaussian primitive counts within the mouth interior between the vanilla training strategies and FTC.
Without Soft Binding. As shown in Tab. 2 and Fig. 7, removing the Soft Binding framework results in a noticeable degradation across all metrics and visibly poorer reconstruction quality. Relying solely on LBS-based deformation substantially weakens the modeling capability, thereby failing to capture fine-grained details and resulting in blurred teeth as well as the absence of smile-induced facial wrinkles.
Without ADC. Removing ADC leads to a significant increase in the LPIPS score, suggesting a degradation in perceptual quality due to the loss of high-frequency details. As shown in Fig. 7, the teeth region and facial wrinkles are insufficiently reconstructed by Gaussians, resulting in overly smooth and less detailed renderings.
Without FPE-AP. Replacing FPE-AP criterion with vanilla positional gradient leads to blurred teeth renderings, as the latter considers only geometric information and fails to add Gaussians in regions with high-frequency details.
Without FTC. For the vanilla training strategy, we set the densification interval based on the number of clusters to ensure that the total number of densification steps matches that of FTC. The teeth region exhibits incomplete reconstruction due to its transiently visible nature.
Without Local Decoder $\phi_{l}$. In this setting, only the global branch is used to model the Gaussian deformation field. As illustrated in Fig. 7, facial wrinkles are reconstructed less accurately, and the teeth region appears blurrier.
Without $\mathcal{L}_{\mathrm{vgg}}$. As shown in Tab. 2, removing the perceptual loss $\mathcal{L}_{\mathrm{vgg}}$ improves SSIM marginally but significantly increases LPIPS score, reflecting overly smooth textures.
Figure 9: Comparison of PSNR performance across different methods during training. STAvatar exhibits the highest training efficiency among all compared approaches.
4.4 Gaussian Number Comparison
To further validate the effectiveness of FTC in transiently visible or frequently occluded regions, we count the number of Gaussian primitives whose UV coordinates fall within the mouth region mask. As illustrated in Fig. 8, we compare the number of Gaussian primitives in the mouth interior with and without FTC on the INSTA dataset. Although some variations exist across identities due to differences in motion and expression within video frames, it is evident that all identities exhibit increased density in the mouth region when using FTC, with an average gain of over 400 primitives (approximately 17%).
4.5 Training Efficiency Analysis
As shown in Fig. 9, we present the training progress and corresponding performance measured in terms of PSNR for different methods. The training epochs follow the settings described in Sec. 4.1 to ensure that all methods fully converge. For MonoGaussianAvatar, only the first 40 epochs are reported because its PSNR shows negligible improvement in the later stages of training. In addition, since different methods adopt distinct initialization strategies, their performance at epoch 0 varies accordingly. Notably, our method, STAvatar, exhibits the highest training efficiency among all compared approaches and almost reaches convergence within 6 epochs, surpassing the second-best method by over 2 dB.
5 Conclusion
In this paper, we present STAvatar, a novel method for high-fidelity and training-efficient reconstruction of animated avatars from monocular videos. The proposed UV-Adaptive Soft Binding framework enables flexible and non-rigid deformation field modeling while remaining compatible with the Adaptive Density Control (ADC), thereby effectively capturing subtle expressions and fine-grained details. Besides, our Temporal ADC strategy addresses the limitations of vanilla ADC in dynamic avatar reconstruction, resulting in more accurate and complete reconstructions in frequently occluded regions. Extensive experiments demonstrate that our method significantly outperforms previous approaches in both reconstruction quality and training efficiency.
Acknowledgement
This work was supported in part by Beijing Natural Science Foundation L242092, Chinese National Natural Science Foundation Projects 92570119, 62276254, U23B2054, the Science and Technology Development Fund of Macau Project 0140/2024/AGJ, and InnoHK program.